OpenStudy (anonymous):

divide. 10c^3/b^3c DIVIDED BY 4c/3b^2

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

\[\frac{10c^3}{b^3c} / \frac{4c}{3b^2} =\frac{10c^3}{b^3c} * \frac{3b^2}{4c} \]

OpenStudy (anonymous):

Now just simplify :)

OpenStudy (anonymous):

mmk so 30c^3b^2/4b^3c^2 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now just simplify this

OpenStudy (anonymous):

30c/4b

OpenStudy (anonymous):

30/4 is 7.5

OpenStudy (anonymous):

that's right, but you can leave as \[\frac{15c}{2b}\]

OpenStudy (anonymous):

thank you =)